# SAGe: A Lightweight Algorithm-Architecture Co-Design for Alleviating Data Preparation Overheads in Large-Scale Genome Analysis

### v1

@@ 4. SAGe: Overview
We propose SAGe, a scheme for storing and acessing large-scale genomic data that not only stores genomic data in a space-efficient compressed way, but also serves requested data in the desired format in a fast and energy-efficient manner.
We address the challenges of optimizing both the performance of preparing data and the space-efficiency of storing data while maintaining low hardware costs via holistic co-design of new (de)compression algorithm, light- weight hardware units, storage data layout, and interface commands, leveraging the features of genomic data. Doing so enables rapid interpretation and decoding via lightweight hardware, while relying on efficient sequential accesses when accessing stored genomic data. [^7] Since compression time is not on the critical-path of genome analysis pipelines, SAGe’s compression and encoding is performed on the host system. When writing the SAGe-compressed data, SAGe leverages its customized interface commands (§5.3) and FTL (§5.4) to efficiently map the data for its efficient access during decompression.
Overview. SAGe supports data preparation for genome analysis systems, regardless of where they are implemented. For example, Fig. 4 shows SAGe’s implementation inside and outside an SSD. The in-storage implementation of SAGe enables in-storage processing for genome analysis in practice (i.e., without relying on large-scale genomic data to be stored in uncompressed formats). On the other hand, SAGe’s implementation outside the SSD can provide benefits when two conditions are met: (i) the given genomic dataset and application do not benefit from ISP, and (ii) the SSD’s external bandwidth limits application performance (e.g., when an in-storage filter cannot efficiently work for the target dataset, and the bandwidth between the SSD and the genome analysis accelerator outside the SSD is limited). In such a case, SAGe’s lightweight hardware can be tightly integrated with the genome analysis accelerator to quickly prepare data and pass it to the accelerator at a high bandwidth.
Fig. 4(a) shows a high-level overview of SAGe’s operations, in the case of its in-storage data preparation. When the host requests genomic data ( 1 in Fig. 4(a)) using SAGe’s interface commands (§5.3), SAGe prepares for its operations based on the SAGe FTL (§5.4) metadata ( 2). The SAGe FTL manages communication between the SSD and the genome analysis systems and coordinates data flow between different SSD parts (e.g., flash chips and in-storage accelerators). After preparation, SAGe accesses genomic data in its SAGe-compressed format (§5.1) from flash chips ( 3) and decompresses it into the desired format using SAGe’s accelerators (§5.2) in the SSD controller ( 4). Based on the interface commands, the decompressed data can then go through genome analysis inside ( 5) and/or outside the SSD ( 6).
Figure 4. Overview of SAGe with its accelerator unit (a) inside and (b) outside the SSD.
Based on our design, SAGe’s data preparation outside the SSD is straightforward: As shown in Fig. 4(b), SAGe’s accelerators need to be implemented outside the SSD and use SAGe-specific I/O interfaces and FTL to manage data layout directly inside the SSD. Due to its lightweight design, SAGe’s hardware units outside the SSD can easily be integrated with the genome analysis systems either directly on the same chip (i.e., with the genome analysis accelerator, which can be a stand-alone ASIC (turakhia2018darwin, ; fujiki2018genax, ; madhavan2014race, ), or implemented in an FPGA (saavedra2020mining, ; zhang2023genomix, ; cervi2022metagenomic, ; shih_haru_2023, ; fujiki2020seedex, ; banerjee2018asap, ; goyal2017ultra, ; chen2016spark, ; chen2014accelerating, ; chen2021high, ; fei2018fpgasw, ; waidyasooriya2015hardware, ; chen2015novel, ; rucci2018swifold, ; haghi2021fpga, ; li2021pipebsw, ; ham2020genesis, ; ham2021accelerating, ; wu2019fpga, ; liyanage2023efficient, ), near memory (wu2021sieve, ; shahroodi2022krakenonmem, ; shahroodi2022demeter, ; dashcam23micro, ; hanhan2022edam, ; zou2022biohd, ; cali2020genasm, ; huangfu2018radar, ; khatamifard2021genvom, ; gupta2019rapid, ; li2021pim, ; angizi2019aligns, ; zokaee2018aligner, ; Zhang_2023_alignerD, ), etc.) or connected via a PCIe interface (i.e., similar to how individual accelerators or GPUs connect to the system).  [^8] In our evaluations, we consider the PCIe-connected configuration for when SAGe’s hardware units are outside the SSD.
For its implementation in or near storage, it is possible for SAGe’s operations to run either on our lightweight specialized accelerators or, alternatively, on other general-purpose in-storage (e.g., (torabzadehkashi2019catalina, ; zou2022assasin, )) or near-storage (e.g., Samsung SmartSSD (lee2020smartssd, )) processing systems. This is because, leveraging our optimizations in algorithms and data structures, SAGe’s operations require only simple operations and small buffers. For example, SAGe’s hardware units consume only 2.5% of the lookup tables and 0.8% of the flip-flop resources of SmartSSD’s FPGA (lee2020smartssd, ). [^9] Since SmartSSD’s FPGA is connected to the SSD with a PCIe lane, it can act as the SGout (SAGe outside the SSD) configuration analyzed in Fig. 12 and provide similar benefits. This emphasizes SAGe’s ease of adoption. Ultimately, choosing between these SAGe configurations (specialized lightweight hardware or general in-/near-storage units) is a design decision with different trade-offs, with the accelerator achieving higher power efficiency. Efficiently leveraging these underlying hardware units relies on SAGe’s optimized data mapping, interface commands, and algorithms.
@@ 5. SAGe: Detailed Design
5.1. Compression and Data Structures
@@
Fig. 5 shows an overview of SAGe’s compression mechanism. SAGe also takes the consensus-based approach widely used in existing genomics-specific compression techniques (chandak2017compression, ; dufort2021renano, ; kokot2022colord, ; kowalski2019pgrc, ); for each read, SAGe stores the mismatch information between the read and a consensus sequence ( 1 in Fig. 5). [^10] Note that, as detailed in §2.3, identifying matches against the consensus during compression does not eliminate need for read mapping during later analyses (Siren2024, ; Sedlazeck2018, ). The key difference of SAGe over existing techniques is how it encodes the mismatch information; instead of using expensive general-purpose compression algorithms (e.g., LZMA (pavlov2015lzma, ), the BWT-based BSC (grebnov2011libbsc, ) or FM-Index search (ferragina2000opportunistic, ), etc.), SAGe (i) stores the information in simple structures that can be efficiently decoded by lightweight hardware with streaming data access, and (ii) minimizes the sizes of these structures by leveraging the features of read sets.
To this end, SAGe sequentially stores the mismatch information (in this figure’s example, delta-encoded mismatch positions) of different reads back-to-back ( 2) in a mismatch position array (MPA). SAGe then improves the capacity efficiency ( 3) by (i) picking the most suitable set of bit count values to use for each entry in the MPA and (ii) using a lightweight structure, i.e., mismatch position guide array (MPGA), to encode which bit count has been used to store each entry in the MPA. In this example, a value of 0 or 1 at index $i$ of the MPGA specifies that the corresponding entry in the MPA is stored using one or four bits, respectively. This is based on the key observation that the distribution of delta-encoded values of mismatch positions is heavily skewed towards smaller values (reasons in §5.1.1). During compression, to minimize the overall data size, SAGe tunes the parameters of each array and guide array ( 4) based on the features of each read set, such as read length and mismatch distribution, which depend on the sequencing technology and the genetic composition of the read set. The configuration parameters of the array and guide array bit count values used for compressing each read set are then encoded at the beginning of the corresponding compressed read set file. When the read set is accessed, SAGe decodes these parameters and utilizes them during the decompression process (§5.2.3).
@@
This data encoding enables decoding with only efficient sequential scans through the data structures. Since SAGe stores mismatch information in the order of appearance in each read, decoding this information is feasible via a sequential scan through the array structures. Since SAGe stores the reads’ positions based on their order of appearance in the consensus sequence, [^11] As discussed in §2.3, various other genomic compressors (e.g., (chandak2017compression, ; chandak2018spring, ; roguski2018fastore, ; cogo2021genodedup, )) also store reads in this order, which enables delta encoding for capacity-efficiency. SAGe also only needs sequential scans through the consensus. Due to this efficient encoding, SAGe can directly send each read to the genome analysis system as soon as it is decoded. Since genome analysis applications can analyze reads independently, they can begin analyzing reads as soon as they are received from SAGe.
By leveraging all of the features that enable SAGe to decode data using only lightweight operations and efficient scans through the data, SAGe is also inherently memory-hierarchy friendly on general-purpose systems due to two reasons. First, as opposed to the state-of-the-art compression techniques used for genomic data that require a large number of random accesses to the main memory, SAGe relies only on efficient streaming accesses. Second, given that all data is accessed sequentially with full spatial locality, each cache line is fully used, which leads to fewer accesses to main memory compared to other algorithms that require many fine-grained random accesses. These factors make SAGe suitable also for cases where the user wants to use it purely in software on general-purpose systems.
5.1.1. Mismatch Positions and Mismatch Counts.
SAGe stores mismatch positions for different reads in a read set sequentially in MPA and uses MPGA to indicate the number of bits used for each MPA entry. SAGe stores the number of mismatches in each read followed by the mismatch positions in that read. This allows SAGe to determine how many elements of MPA and MPGA it needs to access in order to retrieve all mismatch positions for each read. To reduce the sizes of MPA and MPGA, SAGe introduces data-driven optimizations that leverage the distinctive characteristics of short and long read sets. Fig. 6 demonstrates these characteristics for representative real-world genomic read sets (details in §6).
@@
Figure 6. (a) Distribution of #bits needed to store the delta-encoded mismatch positions, (b) Distribution of #reads with different mismatch counts, (c) Cumulative distribution of indel block lengths, (d) Cumulative distribution of #bases in indel blocks of different lengths.
@@
First, Fig. 6(a) shows the distribution of the number of bits required to store the delta-encoded mismatch positions for a representative long read set (RS4 in Table 3). We observe that the distribution is skewed towards the smaller values. This is due to two main factors that can lead to nearby mismatches. First, mutations tend to cluster in some regions of the genome (Bourque2018, ). Second, a local drop in base quality during sequencing can cause incorrect bases in the read (LaPierre2019, ). Based on this observation, SAGe performs two optimizations to minimize the overall size of MPA and MPGA. First, during compression for each read set, SAGe tunes (i) the number of distinct bit counts used for MPA, and (ii) the specific values of each bit count. To do so, SAGe generates the distribution of bits needed to store all mismatch positions in a read set. Based on this distribution and examining various combinations of the bit count values used in MPA and MPGA, SAGe selects the combination that minimizes the total size of MPA and MPGA in the read set. Second, SAGe refines the guide array encoding by exploiting the higher frequency of smaller bit counts. For example, with four different bit counts ($N_{0}$ to $N_{3}$), rather than using 00, 01, 10, and 11 to represent them, SAGe uses 0, 10, 110, and 1110. This encoding can then be efficiently decoded via sequentially scanning MPGA.
@@
Second, we analyze the mismatch counts per read in Fig. 6(b) for a representative human short read set (RS2 in Table 3). As also shown by prior works (e.g., (nag2019gencache, ; mansouri2022genstore, )), most short reads have no or few mismatches. To alleviate the overhead of storing #mismatches for most reads, we use the adaptive encoding mentioned above.
@@
Third, recognizing that indel blocks (consecutive inserted or deleted bases) cause many mismatches in long reads, we analyze their characteristics in a representative read set (RS4 in Table 3). Figures Fig. 6(c) and Fig. 6(d) show the cumulative distributions of indel block lengths and the bases stored for different block lengths, respectively. We observe that (i) most indel blocks are of length one—a common trait in read sets (ono2020pbsim2, ; belyaeva2022best, ), and (ii) although single-base blocks are more frequent, larger blocks encompass most indel bases. Based on these, we implement two optimizations. First, we store the position of the first mismatch and the length of the indel, rather than storing each mismatch position. Second, after detecting an indel (see §5.1.2), we reserve one bit in MPGA to indicate if it is a single-base indel; if not, we allocate 8 bits in MPA for the block length, thus saving space for most single-base indels.
@@
Fig. 7 shows an example of the mismatch position data structures and decompression process for long reads, demonstrating MPGA ( 1), MPA ( 2), and how MPGA entries encode bit counts in MPA ( 3). To retrieve mismatch positions, SAGe first reads the mismatch count for Read#1 in MPGA. It then scans MPGA to obtain the bit counts for each mismatch position and reads the specified number of bits from MPA to decode these positions. For example, if MPGA’s entry for a mismatch position is 10, SAGe reads 4 bits in MPA to find that mismatch position. If a mismatch is an indel (e.g., Read#1 Mismatch#2), SAGe checks MPGA if the indel length is greater than 1. If so, it reads the next eight bits from MPA to determine the indel length. After decoding all three mismatch positions of Read#1, SAGe proceeds to the next read. These lightweight structures enable SAGe to store mismatch positions in a capacity-efficient way and decode them using sequential scans.
@@
SAGe stores information regarding mismatching bases/types in a mismatching base and type array (MBTA). Similar to MPA, MBTA is filled sequentially with information on mismatches in different reads. SAGe includes two optimizations in MBTA . First, in long reads, we observe that a significant fraction of mismatching bases (e.g., up to 80% in our datasets) can originate from chimeric reads, which are reads with sequences joined from different regions of the genome due to sequencing or library preparation errors or structural variations (guan2016structural, ). Parts of these reads map to different locations in the consensus sequence. Fig. 8 shows an example of a chimeric read and its two matching positions. The read has eight mismatches in matching position#1 and nine mismatches in matching position#2. Therefore, only accounting for the best matching position (i.e., matching position#1 in this example), as commonly done in prior works (e.g., (dufort2021renano, ; kowalski2019pgrc, )), results in a large number of mismatches. To avoid this overhead, SAGe considers the top $N$ matching positions [^12] In our analysis, $N=3$ leads to the best compression ratio. This parameter can be tuned seamlessly. for chimeric reads and reconstructs the read based on them. For example, considering both matching positions in Fig. 8 leads to only three mismatches, which is smaller than the eight mismatches found at the best matching position. Since, in real chimeric long reads, the number of mismatches at individual matching positions can be substantial, it is more efficient to reconstruct the read from several matching positions (and store the extra matching positions), rather than storing a large number of mismatch bases at only the top matching position.
@@
Second, since in short reads, substitutions constitute a large fraction of mismatches, we devise an optimization that eliminates the need to store types for substitutions. Instead of using two bits to encode each base (A, C, G, T) and two bits to encode each type (substitution, insertion, deletion), we merge base and type encodings to avoid storing types for substitutions. When reconstructing a read during decompression, SAGe checks whether the mismatching base is the same as the consensus sequence at the corresponding mismatch position. If it is, SAGe recognizes that the mismatch is not a substitution but an indel and uses one bit after the mismatching base to distinguish between an insertion or a deletion.
@@
Fig. 9 shows the distribution of the number of bits needed to store delta-encoded matching positions for a representative real human short read set. We observe that most delta-encoded matching positions can be encoded with very few bits. This is due to its sequencing depth, which results in genomic reads being closely spaced, leading to small differences between consecutive matching positions. Since mismatch positions also exhibit a similar skew towards smaller values (as shown in §5.1.1), we apply the same optimization used for mismatch positions to matching positions as well by introducing the matching position array (MaPA) and matching position guide array (MaPGA).
@@
Although most reads are compressed using the encodings in §5.1.15.1.3, two mismatch types require special handling. First, reads containing unidentified bases (N) expand the DNA alphabet to five characters, making 2-bit encoding impossible. Second, some reads include clips—large insertion blocks at the beginning or end. We classify a read as a corner case if it has a mismatch at position 0. We then use a single bit in MBTA to indicate whether a read has a mismatch at position 0 or is a corner case.
@@
Similar to various other genomic compressors (e.g., (chandak2017compression, ; Meng2023, )), SAGe does not include specialized support for quality scores and, instead, can integrate with any quality score (de)compressor (e.g., (dufort2020enano, ; kokot2022colord, )) outside the SSD. This is because, despite the importance of compression ratios for quality scores, quality score decompression time is not on the critical path of many genome analysis pipelines.
5.2. SAGe’s Hardware Design
SAGe’s accelerator recognizes SAGe’s data structures and quickly decompresses them to provide the application with data in its desired format. Fig. 10 shows SAGe’s hardware units and their integration with the SSD in the case of their in-SSD implementation. We discuss SAGe’s data mapping and fetching from the SSD (§5.2.1), the hardware components’ details (§5.2.2), and their operations (§5.2.3). For its outside-SSD implementation, SAGe places the hardware components outside the SSD but still relies on its efficient data mapping inside the SSD.
5.2.1. Mapping and Fetching Data.
SAGe optimizes data mapping in the SSD and data fetching to the hardware accelerator such that (i) the accelerator leverages SSD’s full bandwidth and (ii) does not require costly hardware resources (e.g., large buffer spaces).
To efficiently leverage all SSD channels, SAGe uniformly partitions the data between the flash channels. This efficient partitioning is enabled by SAGe’s sequential access pattern to all of its data structures. Each partition of the consensus sequence, along with the compressed mismatch information of the reads mapped to that partition, is then placed in a separate channel.
To enable accessing data without requiring large buffer spaces, which is particularly critical for when SAGe’s hardware units need to integrate with genome analysis systems in resource-constraint environments (e.g., near-data processing systems within the SSD), SAGe adopts two optimizations. First, we adopt an approach similar to (zou2022assasin, ) and operate on data fetched from NAND flash chips without needing to buffer them in the SSD’s low-bandwidth internal DRAM. Second, as we no longer rely on DRAM buffers, we must minimize the SRAM buffer sizes necessary for performing computations directly on NAND flash data streams. We do so based on two key features enabled by SAGe’s algorithm and data structures: (i) SAGe’s operations are lightweight and perform only streaming accesses, thus do not need to buffer large amounts of data during computation, and (ii) SAGe employs per-channel accelerator units, and since data can be uniformly partitioned across channels, these units operate independently without requiring synchronization or inter-channel buffering. Leveraging these features, SAGe performs operations on flash data streams using a lightweight double-buffering technique. We use one register to load a chunk of data, and another to hold the subsequent data chunk being read from the flash chips. For this double-buffering, we use two 64-bit registers as they are sufficient to fully utilize our per-channel accelerators.
To leverage the full bandwidth of each channel, We devise a specialized data mapping scheme. To this end, we store different SAGe data structures (i.e., the consensus sequence, guide arrays, and position arrays) sequentially in an interleaved manner, such that SAGe only requires sequential access through data in each channel. When a chunk of data is ready in the 64-bit register, SAGe’s hardware accelerator distributes the arrays and guides arrays to the respective registers in the accelerator’s components (§5.2.2) based on the values indicated by the guide arrays (as explained in §5.1). Details on SAGe FTL’s logical-to-physical mapping in flash blocks across different channels are provided in §5.4.
Figure 10. Overview of SAGe’s hardware units.

### v2

@@ 4. SAGe: Overview
We propose SAGe, an algorithm-architecture co-design for storing and acessing large-scale genomic data that (i) mitigates the data preparation bottleneck, (ii) enables storing data in a highly compressed way, and (iii) is lightweight for seamless integration. SAGe supports data generated with different sequencing technologies, including both short and long reads. We address the challenges of optimizing data preparation performance and compression ratios, at low hardware cost, via a synergistic co-design based on the features of genomic data.
SAGe’s co-design comprises five aspects. First, SAGe sequentially encodes information regarding reads in a genomic read set in data structures 5.1) that can be decoded by efficient accesses and lightweight hardware. We efficiently leverage the long-range similarity common in DNA data to compactly, and yet losslessly, encode genomic reads in these structures. Second, despite relying solely on lightweight structures, SAGe achieves high compression ratios by exploiting features of genomic data to optimize the encoding of these structures 5.1) for each dataset. Leveraging these genomic features is key to maintaining high compression ratios without resorting to complex or resource-intensive techniques. Third, we design hardware units 5.2), which can efficiently interpret and decode SAGe’s data structures with efficient streaming accesses and simple operations, thus not requiring expensive resources such as large buffer spaces or memory bandwidth. This enables SAGe to seamlessly and unobtrusively integrate with genome analysis systems at low hardware cost and without competing for resources required by genome analysis tasks. Fourth, we design an efficient data layout 5.3) in the storage system that enables leveraging the storage system’s full bandwidth when accessing genomic data. Fifth, we propose specialized interface commands 5.4), which are exposed to genome analysis applications to communicate with the storage system and SAGe’s hardware to access data and decompress it to the accelerator’s desired format.
Fig. 4 shows the high-level overview of how SAGe handles data preparation for a genome analysis system. When the genome analysis system requests data using SAGe’s interface commands ( 1 in Fig. 4(a)), SAGe starts operating based on its mapping metadata ( 2), which contains information for SAGe’s efficient data layout. SAGe receives genomic data in its compressed format from the storage device ( 3), decompresses it into the desired format using SAGe’s hardware ( 4), [^7] Since compression time is not on the critical-path of genome analysis pipelines, SAGe’s compression is performed on the host system. When writing the SAGe-compressed data, SAGe leverages its customized interface commands (§5.4) and data mapping (§5.3) to efficiently map the data for efficient access during decompression. and feeds it to the genome analysis system ( 5). SAGe’s lightweight design enables it to efficiently integrate with a broad range of genome analysis systems. We demonstrate case studies of SAGe’s integration with different types of systems in §6.
Figure 4. High-level overview of SAGe.
@@ 5. SAGe: Detailed Design
5.1. Compression Algorithm and Data Structures
@@
Fig. 5 shows an overview of SAGe’s compression. SAGe also takes the consensus-based approach (§2.2) widely used in existing genomic compression techniques (e.g., (chandak2017compression, ; dufort2021renano, ; kokot2022colord, ; kowalski2019pgrc, )); for each read, SAGe stores the mismatch information between the read and a consensus ( 1 in Fig. 5). [^8] Note that, as detailed in §2.2, identifying matches against the consensus during compression does not eliminate the need for read mapping in later analyses (Siren2024, ; Sedlazeck2018, ). The key difference of SAGe’s compression over existing techniques is the encoding of the mismatch information: instead of using expensive compression algorithms, SAGe (i) stores the information in data structures that can be efficiently decoded by lightweight hardware, and (ii) minimizes the structure sizes by leveraging the features of read sets. To this end, SAGe sequentially stores the mismatch information—in this example, delta-encoded mismatch positions—from different reads contiguously ( 2) in an array (here, the Mismatch Position Array, or MPA). SAGe then adapts the bit width of each array entry and uses a guide array (here, Mismatch Position Guide Array, i.e., MPGA) to enable correct decoding ( 3). This is based on the observation that the distribution of delta-encoded values of mismatch positions is heavily skewed towards smaller values (reasons in §5.1.1). To achieve high compression ratios using only these lightweight structures, during compression, SAGe tunes the parameters of each array and guide array, e.g., the number of distinct bit-count values used, based on the features of each read set ( 4), as described in §5.1.1, 5.1.2, and 5.1.3. These features depend on the sequencing technology and the genetic composition of the read set. The tuned parameters are then encoded at the beginning of the compressed read set file.
@@
This encoding enables decompression with only sequential accesses through the data structures and does not rely on expensive resources (e.g., large buffers). Since SAGe stores mismatch information in the order of appearance in each read, decoding this information is feasible via sequential accesses through the array structures. Similarly, since SAGe stores the reads’ positions based on their order of appearance in the consensus sequence, [^9] As discussed in §2.2, various other genomic compressors (e.g., (chandak2017compression, ; chandak2018spring, ; roguski2018fastore, ; cogo2021genodedup, )) also store reads in this order, which enables delta encoding for capacity-efficiency. SAGe only needs sequential accesses through the consensus. SAGe can directly send each read to the genome analysis system as soon as it is decoded. Since genomic applications can analyze reads independently, they can begin analyzing reads as soon as they are received from SAGe.
5.1.1. Mismatch Positions and Counts.
@@
Figure 6. Distribution of (a) #bits needed to store the delta-encoded mismatch positions, (b) #reads with different mismatch counts. Cumulative distribution of (c) indel block lengths, (d) #bases in indel blocks of different lengths.
@@
SAGe stores mismatch positions for different reads in a read set sequentially in MPA and uses MPGA to indicate the number of bits used for each MPA entry and the number of mismatches in each read. To reduce the sizes of MPA and MPGA, SAGe introduces optimizations that leverage the features of short and long read sets. Fig. 6 demonstrates these characteristics for representative real-world genomic read sets (details in §7). Fig. 6(a) shows the distribution of the number of bits needed to store the delta-encoded mismatch positions for a representative long read set (RS4 in Table 3). We observe that the distribution is skewed towards the smaller values. This is due to two main factors that can lead to nearby mismatches. First, genetic mutations tend to cluster in some regions of the genome (Bourque2018, ). Second, a local drop in sequencing quality can cause incorrect bases in the read (LaPierre2019, ). Based on this observation, SAGe performs two optimizations. First, during compression of each read set, SAGe tunes (i) the number of distinct bit counts used for MPA, and (ii) the specific values of each bit count. To do so, SAGe generates the distribution of bits needed to store all mismatch positions in the read set. Based on this distribution and examining various combinations of the bit count values in MPA and MPGA, SAGe selects the combination that minimizes total size. Second, SAGe exploits the higher frequency of smaller bit counts to refine the guide array encoding by assigning shorter representations to more common inputs. For example, assuming four distinct bit counts, SAGe uses 0, 10, 110, and 1110 to represent them, rather than using 00, 01, 10, and 11.
@@
Fig. 6(b) shows mismatch counts per read for a representative human short read set (RS2 in Table 3). As also shown by prior works (e.g., (nag2019gencache, )), most short reads have no or few mismatches. To alleviate the overhead of the frequent occurrence of low mismatch counts, we use the adaptive encoding mentioned above.
@@
Recognizing that indel blocks (consecutive inserted or deleted bases) cause many mismatches in long reads, we analyze their characteristics in a representative read set (RS4 in Table 3). Figures Fig. 6(c) and Fig. 6(d) show the cumulative distributions of indel block lengths and the bases stored for different block lengths, respectively. We observe that (i) most indel blocks are of length one—a common trait in read sets (ono2020pbsim2, ; belyaeva2022best, ), and (ii) although single-base blocks are more frequent, larger blocks encompass most indel bases. Based on these, we implement two optimizations. First, we store the position of the first mismatch and the length of the indel, rather than storing each mismatch position. Second, after detecting an indel (see §5.1.2), we reserve one bit in MPGA to indicate whether it is a single-base indel. If not, we allocate eight bits in MPA to store the block length, thus avoiding space waste for the majority of indel blocks, which are of length one.
@@
Fig. 7 shows an example of the decompression process for mismatch positions in long reads, demonstrating MPGA ( 1), MPA ( 2), and how MPGA entries encode bit counts in MPA ( 3). SAGe first reads the mismatch count for Read#1 in MPGA. It then scans MPGA to obtain the bit counts for each mismatch position and scans the specified number of bits from MPA to decode these positions. For example, if MPGA’s entry for a mismatch position is 10, SAGe reads 4 bits in MPA to find that mismatch position. If a mismatch is an indel (e.g., Read#1 Mismatch#2), SAGe checks MPGA if the indel length is greater than 1. If so, it reads the next eight bits from MPA to determine the indel length. After decoding all three mismatch positions of Read#1, SAGe proceeds to the next read.
@@
SAGe stores information regarding mismatching bases/types in a Mismatching Base and Type Array (MBTA). Similar to MPA, MBTA is filled sequentially with information for different reads. SAGe includes two optimizations based on the features commonly seen in genomic data. First, in long reads, we observe that a significant fraction of mismatching bases (e.g., up to 80% in our datasets) can originate from chimeric reads, which are reads with sequences joined from different regions of the genome due to sequencing or library preparation errors, or structural variations (guan2016structural, ). Parts of these reads map to different locations in the consensus. Fig. 8 shows a chimeric read with eight mismatches at matching position#1 and nine at position#2. Thus, considering only the best matching position, as in prior works (e.g., (dufort2021renano, ; kowalski2019pgrc, )), results in many mismatches. Instead of relying on expensive compression algorithms to compress these mismatches, SAGe considers the top $N$ matching positions for chimeric reads. For example, considering both positions in Fig. 8 reduces mismatches to three, fewer than the eight at the best single position. Since, in real chimeric reads, the number of mismatches at individual positions can be substantial, reconstructing reads from multiple positions (and storing these positions) is more efficient than storing many mismatches from only the top position.
@@
Second, since substitutions dominate mismatch types in short reads, instead of using two bits for each base (A, C, G, T) and two bits for each type (substitution, insertion, deletion), we merge their encodings to avoid storing types for substitutions. When reconstructing a read, SAGe checks whether the mismatching base is the same as the consensus at the corresponding mismatch position. If so, SAGe recognizes that the mismatch is not a substitution but an indel and uses one bit after the mismatching base to distinguish between an insertion or a deletion.
@@
Fig. 9 shows the distribution of the number of bits needed to store delta-encoded Matching positions for a representative real human short read set. We observe that most delta-encoded matching positions can be encoded with very few bits. This is due to its sequencing depth, which results in genomic reads being closely spaced, leading to small differences between consecutive matching positions. Since mismatch positions also exhibit a similar skew towards smaller values (as shown in §5.1.1), we apply the same optimization used for mismatch positions by introducing the matching position array (MaPA) and Matching Position Guide Array (MaPGA). SAGe then tunes the array encodings based on the matching position distribution for each read set.
@@
Two corner cases require special handling. First, reads containing unidentified bases (N) expand the DNA alphabet to five characters, making 2-bit encoding impossible. Second, some reads include clips (i.e., large insertion blocks at the beginning or end) that need reattachment during decompression. Adding indicator bits to specify these corner cases would significantly increase overhead (particularly for short reads). To avoid this, we classify a read as a corner case if it has a mismatch at position 0. We then use a single bit in MBTA to indicate whether a read has indeed a mismatch at position 0 or is a corner case, thus efficiently identifying these cases without impacting standard reads.
@@
Like various other genomic compressors (e.g., (chandak2017compression, ; Meng2023, )), SAGe does not specialize for quality scores and, instead, can integrate with any quality score (de)compressor (e.g., (dufort2020enano, ; kokot2022colord, )). This is because quality score decompression is not on the critical path of many genome analysis pipelines for two reasons. First, they are accessed in later stages of analysis pipelines, i.e., for variant calling after read mapping (Poplin2018, ; li2008mapping, ; li2009sequence, ). Second, variant calling only requires quality scores for the small fraction of positions surrounding mismatches (Poplin2018, ; li2008mapping, ).
5.2. Hardware for Data Preparation
Figure 10. Overview of SAGe’s hardware.